A disc of radius R is kept such that its axis coincides with the x-axis and its centre is at (d,0,0) .The thickness of disc is t and it carries a uniform volume charge density ¶ .The external electric field in the space is given by E=kr(r is position vector of any point in space with respect to the origin of the coordinate system). If the electric force on the disc is mπkdt¶R^2 /4 then find value of m?

Grade:12

1 Answers

Solitary
3 years ago
Given:
radius of disc = R
thickness = t
volume charge density = \rho
it’s placed at (d,0,0) with its axis along the x-axis
 
since the thickness is t, the surface charge density of each surface can be given as
\sigma = \rho t
now on the surface, consider a small ring element dr at a distance r from the centre
the area covered by this ring can be given by 2\pi rdr
thus the charge
dq = \sigma\cdot2\pi r dr = \rho t \cdot 2\pi r dr
now, the force can be given as
dF = E\cdot dq
dF = Kd \cdot 2\pi \rho t rdr
integrating both sides,
\int_0^F dF = \int_0^R Kd \cdot 2\pi \rho t rdr
F = 2 \pi Kdt\rho [\frac{r^2}{2}]_0^R
F = \pi K dt \rho R^2 
comparing this with F = m\pi K dt \rho R^2
m = 1
 
Hence, m=1 is the required solution.